A uniform electric field of magnitude 211 V/m is directed in the positive x direction. A -13.5 μC charge moves from the origin to the point (x, y) = (21.7 cm, 54.1 cm). What was the change in the potential energy of this charge? Through what potential difference did the charge move?
Answer #1

Given,

V = 211 V/m ; q = -13.5 uC ; P = 21.7 cm , 54.1 cm

We know that

U = q V

V = Ed

U = q E d

d = x = 21.7 cm = 0.217 m

U = 13.5 x 10^-6 x 211 x 0.217 = 6.2 x 10^-4 J

Hence, U = 6.2 x 10^-4 J

V = U/q

V = -6.2 x 10^-4/(-13.5 x 10^-6) = -45.93 V

Hence, V = 45.93 V